OpenStudy (candycanegone):

A company invested $29,000, putting part of it into a savings account that earned 3.2% annual simple interest and the remainder in a stock fund that earned 12.4% annual simple interest. If the investments earned $1572 annually, how much was in each account?

OpenStudy (wolf1728):

.032x + .124y = 1,572 x + y = 29,000 Then solve for x and y

OpenStudy (candycanegone):

um

OpenStudy (wolf1728):

a) .032x + .124y = 1,572 b) x + y = 29,000 Divide equation "a" by -.032 a) -x -3.875y = -49,125 b) x + y = 29,000 Now add equations "a" and "b" The "x" will zero out and you can then solve for "y"